Circular No. 8846

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1281 SCORPII R. W. Russell, R. J. Rudy, D. K. Lynch, S. Mazuk, and R. L. Pearson,

The Aerospace Corporation; C. E. Woodward, University of Minnesota; R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and R. B. Perry, Langley Research Center, NASA, report 0.4- to 2.5-μm spectroscopy of this nova (cf. IAUC 8810, 8812) using the Lick 3-m telescope (+ VNIRIS) on May 6 UT, as well as 0.8- to 5.5-μm spectroscopy using the Infrared Telescope Facility (IRTF) 3-m telescope (+ SpeX) on May 31. This object has faded considerably over the past month, but the character of its rather complex spectrum has not changed significantly. The spectrum has broad and clearly doubled lines and is dominated by He I emission features. O I is still present, and the forbidden lines [O II], [O III], and [N II] are present. There are no indications of coronal lines or dust formation. The interstellar reddening derived from the O I lines is E(B − V) = 0.7.

V2615 OPHIUCHI Rudy, Lynch, Mazuk, Russell, Pearson, Woodward, Puetter, and Perry,

also report 0.4- to 2.5-μm spectroscopy of V2615 Oph (cf. IAUC 8824), obtained at Lick as above on May 7 UT, and 0.8- to 5.5-μm spectroscopy, obtained at the IRTF as above on May 31. This nova evolved dramatically between the two observations due to the formation of dust. The reddening derived from the O I lines jumped from E(B − V) = 1.0 to 1.3. The emission-line spectrum remained very rich and of low excitation with numerous features of Fe II and N I. The infrared Ca II triplet was present, and the C I lines were unusually strong, even for a nova. The carbon-monoxide emission reported by Das et al. on CBET 925 was no longer present.

COMETS C/2007 D4, C/2007 D5, C/2007 E4, C/2007 E5 (SOHO) Four additional Kreutz-sungrazing comets have been found on SOHO

website images (cf. IAUC 8845) — the “discovery” data tabulated below — all being very faint and stellar in appearance.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 D4 Feb. 16.613 22^{h}19.^{m}3 −13^{o}43′ C3 HS 2007-K66
C/2007 D5  18.821 22 24.0 −12 36 C3 AK 2007-K68
C/2007 E4 Mar. 2.904 23 09.7 − 7 55 C3 BZ 2007-K66
C/2007 E5  9.721 23 33.8 − 5 10 C3 VB 2007-K68
2007 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green